F(x)=x+tanx,f(x) is inverse of g(x) differentiation of g(x) is equal to 

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
Since g is inverse of f, therefore
f(g(x))= x
f'(g(x))*g'(x)=1
But f'(x)= 1+sec^2x
So g'(x)= 1/(1+sec^2g(x))